### v2

# AnyRIR: Robust Non-intrusive Room Impulse Response Estimation in the Wild

Abstract
We address the problem of estimating room impulse responses (RIRs) in noisy, uncontrolled environments where non-stationary sounds such as speech or footsteps corrupt conventional deconvolution. We propose AnyRIR, a non-intrusive method that uses music as the excitation signal instead of a dedicated test signal, and formulate RIR estimation as an $\ell_{1}$-norm regression in the time–frequency domain. Solved efficiently with Iterative Reweighted Least Squares (IRLS) and Least-Squares Minimal Residual (LSMR) methods, this approach exploits the sparsity of non-stationary noise to suppress its influence. Experiments on simulated and measured data show that AnyRIR outperforms $\ell_{2}$-based and frequency-domain deconvolution, under in-the-wild noisy scenarios and codec mismatch, enabling robust RIR estimation for AR/VR and related applications.
Index Terms—  Acoustic measurements, deconvolution, music.
1 Introduction
A room impulse response (RIR) characterizes the acoustic properties of an environment, capturing how sound propagates and reflects within a space. Given a specific RIR, it is possible to render arbitrary sound sources as if they were played within that particular environment [11]. This capability is especially valuable for applications in augmented and virtual reality (AR/VR) and smart speakers, where realistic simulation of various acoustic spaces, from domestic spaces to large public venues, is essential [9, 10].
However, estimating RIRs in uncontrolled real-world environments remains a challenge [11]. Conventional methods, such as exponential sine sweeps, require controlled and often intrusive setups that are impractical in public or dynamic spaces. Moreover, non-stationary noise sources (e.g., footsteps or speech) corrupt these measurements [5, 14, 15]. Inspired by the widespread presence of background music in public venues and the availability of clean references through music identification algorithms such as Shazam [shazam2025, wang_industrial-strength_2003], we propose a robust and non-intrusive method for estimating RIRs in the wild using music as the excitation signal. Fig. 1 illustrates how our method can estimate an RIR from music recorded in a cafe.
Fig. 1: (Top) A noisy environment, such as cafe, where music is playing and is reverberating, and (bottom) usage of AnyRIR to estimate the RIR with the help of existing music recognition algorithm.
While stationary noise with a Gaussian distribution can often be handled adequately by $\ell_{2}$-based deconvolution methods, non-stationary noise violates the underlying assumptions and severely degrades estimation performance [17]. MOSAIC [15] introduces a way to remove non-stationary noise from exponential sine sweep measurements using median filtering, but it requires at least three repeated measurements with exactly the same excitation signal, which makes it somewhat inflexible. As an alternative, we adopt a regression-based approach using the $\ell_{1}$-norm loss, which is known for its robustness to outliers [Boyd:2004jj]. This makes it particularly suitable for handling non-stationary noise, since it suppresses the influence of transient disturbances in the measurement. Although $\ell_{1}$-norm regularization has previously been used to promote sparsity in RIR estimation [crocco2015room], inducing sparsity in the filter is not our objective. Instead, the $\ell_{1}$-norm is used solely as the data fidelity term to improve robustness to non-stationary noise, without imposing sparsity or other constraints on the RIR.
Fig. 2: The proposed AnyRIR method estimates the RIR by minimizing the time-frequency residual between the recorded music and the clean music convolved with the previous RIR estimate. The main tools used are the IRLS and LSMR methods.
Robust deconvolution using the $\ell_{1}$-norm has been explored in other domains. In underwater acoustics, it has been applied to channel estimation under impulsive noise using time-domain ADMM [2]. In image processing, $\ell_{1}$-norm minimization has been widely used to suppress impulsive noise during deconvolution [1, 4]. In contrast, our approach operates in the time–frequency domain and does not impose sparsity on the filter, focusing instead on robustness to non-stationary noise in real-world acoustic measurements.
2 Problem Formulation
Given an RIR $\bm{h}$ and an excitation signal $\bm{x}$, the measured signal $\bm{y}$ at a microphone can be modeled as $\bm{y}=\bm{x}\mathbin{\ast}\bm{h}+\bm{n},$ where $\mathbin{\ast}$ denotes the convolution operation and $\bm{n}$ represents additive background noise. Let $\bm{x}$ have length $L$, and $\bm{h}$ have length $N$, then $\bm{y}$ has length M=L+N1M=L+N-1. We aim to solve the inverse problem of estimating the unknown RIR $\bm{h}$ from a known pair of excitation and measurement signals, $\bm{x}$ and $\bm{y}$. When the noise $\bm{n}$ is assumed to be Gaussian, modeling stationary background noise, the deconvolution problem can be approached with the $\ell_{2}$-norm loss. However, real-world noise often contains non-stationary impulsive components, which is better modeled by heavy-tailed, non-Gaussian distributions. Such impulsive components act as outliers in the residual and degrade the performance of $\ell_{2}$-based estimators. To enhance robustness against outliers, we adopt the $\ell_{1}$-norm loss [Boyd:2004jj] and write the convolution as a multiplication operator by forming the Toeplitz matrix $\bm{X}$ of $\bm{x}(t)$:
$\begin{aligned} \hat{\bm{h}}&=\operatorname*{arg\,min}_{\bm{h}}\norm{\bm{y}-\bm{x}\mathbin{\ast}\bm{h}}_{1}=\operatorname*{arg\,min}_{\bm{h}}\norm{\bm{y}-\bm{X}\bm{h}}_{1}.\end{aligned}$ (1)
Furthermore, many non-stationary disturbances are sparse in the time-frequency domain, e.g., speech [16, 13]. Thus, we perform a time-frequency residual computation:
$\begin{aligned} \hat{\bm{h}}=\operatorname*{arg\,min}_{\bm{h}}\norm{\bm{S}\bm{y}-\bm{S}\bm{X}\bm{h}}_{1},\end{aligned}$ (2)
where $\bm{S}$ is the time-frequency transform matrix, e.g., the stacked matrix of DFT matrices which forms the short-time Fourier transform (STFT) matrix.
3 Proposed Estimation Method
The proposed method, AnyRIR, solves the deconvolution problem using the $\ell_{1}$-norm minimization (2) in the time-frequency domain, where $\bm{X}$ and $\bm{S}$ are large matrices. This leads to a time and memory limitation when using standard optimization libraries, such as CVXPY [diamond2016cvxpy]. There are various efficient alternatives, such as ADMM [2]. In particular, we choose the Iterative Reweighted Least Squares (IRLS) method along with a matrix-free iterative method called the Least-Squares Minimal Residual (LSMR) to handle the least-squares subproblem [7], as shown in Fig. 2, while avoiding to form $\bm{X}$ and $\bm{S}$ explicitly. The IRLS algorithm is formulated as a sequence of weighted least squares problems:
$\begin{aligned} \bm{h}^{(i+1)}&=\operatorname*{arg\,min}_{\bm{h}}\norm{\bm{w}^{(i)}\mathbin{\circ}\left(\bm{S}(\bm{y}-\bm{X}\bm{h}^{(i)})\right)}_{2},\\ \bm{r}^{(i+1)}&=\bm{S}(\bm{y}-\bm{X}\bm{h}^{(i+1)}),\\ \bm{w}^{(i+1)}&=\frac{1}{\max(\lvert(\bm{r}^{(i+1)}\rvert,\delta^{(i)})},\end{aligned}$ (3)
[…]
For small-scale problems where $\bm{X}$ is well-conditioned, the least squares subproblem can be solved analytically using the normal equations. However, in our setting, where the excitation signal is very long, forming the normal equation becomes memory-intensive. In addition, the convolutional structure of $\bm{X}$ often leads to ill-conditioned systems [3], resulting in numerical instability. Therefore, we employ an iterative solver, such as the Conjugate Gradient method [Shewchuk.1994] and LSQR [12]. In particular, we use LSMR, which is a matrix-free method for sparse and or fast linear operator, offering improved robustness when handling ill-conditioned matrices [7].
LSMR requires both a forward and an adjoint operator, as it does not compute the normal equation explicitly. The forward operator computes $\bm{y}=\bm{S}\bm{X}\bm{h}$, a matrix-form of convolution operator followed by a time-frequency transform. The adjoint operator reverses this sequence, $\bm{h}=\bm{X}^{H}\bm{S}^{H}\bm{y}$, where $\bm{S}^{H}$ is the adjoint of the time-frequency transform and $\bm{X}^{H}$ is the adjoint of convolution, corresponding to the time-reversed excitation signal.
Both convolution and its adjoint are implemented using FFT-based operations, avoiding explicit construction of the convolution matrix $\bm{X}$. The FFT-based convolution and the adjoint convolution are computed as
$\begin{aligned} \bm{X}\bm{h}&=\textrm{trunc}_{N}(\textrm{iFFT}_{M}(\textrm{FFT}_{M}(\bm{x})\mathbin{\circ}\textrm{FFT}_{M}(\bm{h})))\\ \bm{X}^{\mathstrut\scriptscriptstyle H}\bm{y}&=\textrm{trunc}_{N}(\textrm{iFFT}_{M}(\textrm{FFT}_{M}(\textrm{flip}(\bm{x}))\mathbin{\circ}\textrm{FFT}_{M}(\bm{y}))),\end{aligned}$ (4)
where $M$ is the FFT length and $N$ is the output length. Time-frequency transform $\bm{S}$ and its Hermitian $\bm{S}^{H}$ are implemented with the SciPy STFT and iSTFT functions, which are linear, properly paired, and normalized such that the iSTFT acts as the adjoint of the STFT. To achieve this, we apply a boxcar window with no overlap, use zero-padding of $2N_{\mathrm{DFT}}$, and apply a normalization factor of $\sqrt{N_{\mathrm{DFT}}/2}$, since the scipy STFT uses an unnormalized FFT. We fix the DFT length to $N_{\mathrm{DFT}}=256$.
The conditioning of the system matrix $\bm{X}$ is critical for the convergence of iterative solvers in least squares problems. In our case, the system matrix $\bm{X}$ is a Toeplitz matrix. For such a matrix to be well-conditioned, the power spectral density (PSD) should be approximately flat, i.e., the excitation signal should be spectrally white. However, music signals typically do not have a flat spectrum. While preconditioning techniques for such Toeplitz matrices are well studied [3], flattening the PSD of the excitation signal achieves a similar effect. We therefore apply preprocessing to equalize (EQ) both the excitation and measurement signals, as shown in Fig. 2. Specifically, we downsample the audio to 32 kHz, inject high-frequency noise, and apply a shared high-order inverse linear prediction filter (order 200). This step acts as a preconditioner, improving numerical stability and convergence of IRLS-based deconvolution. Fig. 3 presents the improvement in convergence, showing that the EQ preconditioning reduces the number of necessary iterations by a factor of 10. The code for an implementation of the full method is available online [^1] https://github.com/kyungyunlee/robust-deconv.
Fig. 3: Number of LSMR iterations required for convergence with and without EQ preconditioning.
4 Evaluation
We evaluate the proposed method in both simulated and real-world scenarios that reflect typical use cases of AnyRIR, such as in cafés. Our approach is compared against two baselines: an $\ell_{2}$-norm minimizing method in the time-frequency domain, which assumes only Gaussian noise, and a frequency-domain deconvolution method [6]. Furthermore, we assess the robustness of AnyRIR under codec mismatch conditions, where the music played in the environment differs in encoding from the reference identified by the music recognition algorithm, a situation that may frequently occur in practice.
4.1 Dataset
To simulate music played in a typical public environment, we convolve music signals with RIRs and add both stationary and non-stationary noise. Specifically, we generated four different songs using Suno AI[suno2025], covering styles characteristic of café environments: pop, melodic acoustic instrumentation, coffee-shop ambiance, male and female vocals, light percussion, and full-band arrangements. For non-stationary noise, we use the AID dataset [8], which contains anechoic recordings of real-world interferers such as footsteps or coughs. RIRs are from the MIT Acoustical Reverberation Scene Statistics Survey [18], which contains 271 RIR representing everyday acoustic environments.
4.2 Presence of Non-stationary Noise
Fig. 4: (Left) Music signal convolved with an RIR with added stationary and non-stationary noise: $\bm{y}$. (Center) Original non-stationary noise. (Right) Weights $\bm{w}$ computed by AnyRIR to suppress the bins containing the non-stationary noise.
Fig. 5: EDCs from different deconvolution methods compared to the ground truth, showing the superiority of AnyRIR.
We evaluate the robustness of AnyRIR on the task of deconvolution in the presence of non-stationary noise. For this experiment, we generated 50 synthetic examples. Each example consists of a 30-s music excerpt convolved with an RIR, with non-stationary noise events (coughing, cutlery, footsteps, glass jar) inserted at random times without overlap. The total duration of non-stationary noise per sample is between 20–50% of the music signal. Further, stationary noise is added to achieve a fairly high SNR of 50 dB, since stationary background-noise suppression is outside the scope of this work.
The method is designed to treat time–frequency bins containing non-stationary noise as outliers and down-weight them during RIR estimation. Fig. 4 shows an example signal $\bm{x}$ alongside the detected non-stationary noise and the corresponding weights $\bm{w}$ assigned by AnyRIR. As seen by comparing the middle and right panels, the method successfully identified noise-dominated bins and suppressed them with weights close to zero, while preserving the clean signal and stationary background noise with weights near one. This behavior reflects the design of our weighting scheme, which attenuates large deviations while leaving small fluctuations unaffected. The effect on RIR estimation can be seen in Fig. 5, which compares the corresponding energy decay curves (EDCs). Relative to the ground-truth RIR, AnyRIR’s estimate deviates only slightly due to background noise in the late tail. In contrast, the $\ell_{2}$-norm and frequency-domain deconvolution methods produce substantially poorer estimates.
As an upper bound, we consider performance on signals without non-stationary noise. Stationary noise remains challenging to suppress, but its effect can be averaged out by extending measurement signals, e.g., to 2–3 min. Table 1 shows that AnyRIR performs equivalently to the standard $\ell_{2}$ method in the absence of non-stationary noise, while providing robustness when such noise is present. Results are reported as mean and standard deviation across the 50 examples.
Table 1: Comparison of RIR estimation error ($\bm{h}$ error in dB, mean $\pm$ std) under different noise conditions. AnyRIR maintains low error in both cases, while baseline methods degrade substantially in the presence of non-stationary noise.
AnyRIR L2 Freq. Deconv.
Stat. noise only $42.0\pm 4.8$ $41.7\pm 4.8$ $7.6\pm 4.9$
Stat. + non-stat. $36.0\pm 5.0$ $10.6\pm 6.8$ $2.8\pm 4.5$
4.3 Music Codec
In real-world scenarios, the music played through a loudspeaker may use a different codec than the reference track retrieved by a music recognition algorithm, e.g., one in WAV, the other in MP3. We call this a codec mismatch scenario, which effectively introduces additional background noise.
To simulate this condition, we degraded signals using FFmpeg by encoding and decoding with different codecs. We considered two cases: in the codec mismatch condition, the measurement signal was compressed with MP3 at 173 kbps while the excitation signal was compressed at 64 kbps; in the same codec condition, both signals were taken from the same MP3 file. This setup isolates the effect of codec differences during system identification.
Figure 6 shows that codec mismatches primarily affect the late tail of the estimated RIR, where increased noise is observed. Quantitatively, the mismatch condition leads to approximately 15 dB higher error compared to the same-codec condition (22-22 dB vs. 37-37 dB $\bm{h}$ error). This degradation is consistent with the interpretation of codec mismatch as an effective background noise source.
Fig. 6: Effect of codec mismatch. When the measurement and excitation signals use different codecs, the estimated RIR exhibits increased noise in the late tail, resulting in $\sim$15 dB higher error compared to the same-codec condition.
4.4 Real-world Evaluation
To evaluate AnyRIR under real-world conditions, we conducted measurements in the kitchen space of the Aalto Acoustics Lab. The ground-truth RIR was obtained using the exponential sine sweep method. For evaluation, we played a 1-min AI-generated music excerpt through a Genelec 8020A loudspeaker and recorded the response with an AKG C414 XLS microphone.
During playback, lab members casually generated everyday noises such as door slams, cutlery sounds, and chair movements. We collected three recordings with varying noise densities, two of which included speech. Fig. 7 shows the EDCs of one of the recordings, where AnyRIR provides an estimate close to the ground truth despite the presence of interfering noise. The related audio examples, video, and additional results are available online [^2] https://kyungyunlee.github.io/anyRIR-demo.
Fig. 7: EDC estimated from a real-world recording.
5 Conclusions
This paper introduced AnyRIR, a robust method for estimating RIRs from music as a non-intrusive alternative to exponential sine sweeps. To handle long excitation signals and non-stationary noise, we cast the problem as an $\ell_{1}$-norm optimization, solved efficiently with IRLS and LSMR in the time–frequency domain and further accelerated using FFT-based convolution. Experiments on 50 simulated recordings and three real-world measurements showed that AnyRIR yields accurate estimates and outperforms $\ell_{2}$-norm and frequency-domain deconvolution in the presence of non-stationary noise and codec mismatch.